Solve the system by substitution. Help me please.
x + y = 12
y = 2x

Answer by ptaylor(2198)


x + y = 12---------------eq1
y = 2x ----------------------eq2
Lets substitute y=2x from eq2 into eq1 and we get:
x+2x=12 collect like terms
3x=12 divide both sides by 3
x=4
Now we can substitute x=4 into eq2 and we get:
y=2*4=8
So our answer is:
x=4 and
y=8
CK
From eq1:
x+y=12
4+8=12
12=12
From eq2:
y=2x
8=2*4
8=8

Hope this helps---ptaylor